what is the answer and soultion to- calculate the diamter of a solid cylinder which has a height of 66cm and total surface area of 3m^2
Found 2 solutions by nyc_function, stanbon:
Answer by nyc_function(2741)

Use the following formula:
Surface Area = (2 • π • rē) + (2 • π • r • height)
Let h = 66, π = 3.14
Let surface area = 3m^2
Find the radius first. Then after finding the radius, multiply it by 2 and you will have your diameter.

Answer by stanbon(75887)

what is the answer and solution to- calculate the diamter of a solid cylinder which has a height of 66cm and total surface area of 3m^2
--------------------
Total surface = 2(surface of base) + (circumference of base)(height)
-------
3 = 2((d/2)^2*pi) + (pi*d(66))
3 = (d^2/2)pi + 66pi(d)
Multiply thru by 2 to get:
6 = pi(d^2) + 66pi(d)
--------------
This is a quadratic in "d"
pi(d^2) + 66pi(d) - 3 = 0
---
Use the quadratic formula to solve for "d":
d = [-66pi +- sqrt((66pi)^2-4pi*-3)]/(2pi)
d = [-66pi +- sqrt(43029.7)]/(2pi)
---
Positive solution:
d = 0.0909/(2pi)
d = 0.0145 cm
==================
Cheers,
Stan H.